OpenStudy (anonymous):

Does the matrix 1 , 2 have an inverse? -2 , 4

OpenStudy (anonymous):

so long as the determinant is not zero, it does

OpenStudy (anonymous):

calculate the mod of the matrix

OpenStudy (anonymous):

the determinant of your matrix is \(1\times 4-2\times (-2)=8\) so you should be in good shape